# From Interaction to Impact: Towards Safer AI Agents Through Understanding and Evaluating UI Operation Impacts

### v1

@@ 5.2. Method
To input the UI elements from the UI action traces into the evaluation prompts, we detect UI elements from each screen (Zhang et al., 2021) and format them as an HTML string (Wang et al., 2022). For MLLMs (multimodal LLMs), we use the raw screenshot data. We used 5 different LLMs and MLLMs for the evaluation: (1) GPT4, (Text Only), (2) GPT4 (MLLM), (3) Gemini 1.5 Flash (Text Only), (4) MM1.5 (MLLM) (Zhang et al., 2024), and (5) Ferret-UI (MLLM) (You et al., 2023), an MLLM trained on UI understanding tasks.
@@ 6. Discussion
In this section, we reflect on the research process and results of two studies. We discuss three major points of limitations and future work: (1) the granularity of impact, (2) brainstorming vs. realistically performing actions, and (3) potential usage of the taxonomy.
@@ 6.2. Future Work
Individuals have different perceived feelings over “impacts.” Our taxonomy can act as a reasoning tool to guide AI agents in determining the level of impact and, ultimately, shaping how they respond to those impacts. By using this taxonomy, future AI agents could make their decision-making process more transparent and understandable, showing how they were calibrated to assess the effects of UI actions. Moreover, users could personalize their own policies based on these taxonomy categories. For instance, if an AI agent determines that a UI action requires “multiple steps in a timely manner”, users could customize the outcome. They might set the decision to “minimum impact” if the action does not affect other users, or to “significant impact” if it changes how others perceive them. This potential of the taxonomy brings new possibilities to make AI agents’ decisions more accessible, customizable, and reliable by understanding different aspects of “impacts.”
Base for future UI agent safety research.
Our taxonomy lays the groundwork for understanding post-action impacts on UIs. As noted in the limitations, further research is needed to refine the granularity of impact assessments. Currently, our data annotations classify impacts into three categories: minimal impact (AI can proceed safely), moderate impact (AI should seek user confirmation), and significant impact (AI should halt and defer to human judgment). However, real-world scenarios are often more complex, with no one-size-fits-all policy for AI behavior. These policies are highly dependent on the specific domain and application. We encourage future researchers to build on our work by exploring the relationship between UI action impacts and the policies that AI agents should follow when navigating UIs.
@@ 7. Conclusion
This work aimed at understanding and evaluating the consequences of UI actions by introducing a comprehensive taxonomy of UI action impacts. Through collaborative workshops with domain experts and an extensive data synthesis study, we have captured a diverse range of UI actions that go beyond harmless browsing, focusing instead on actions with potential interactions with other parties that might need intervention. Our findings demonstrate that current datasets do not adequately represent the complexity of UI action interactions, particularly those with significant consequences. We show a significant gap in LLM’s ability to reliably understand the complexity of what happens as impact after UI actions.
@@ A.4. Adding Chain-of-Thought (CoT)
Roll Back Effects: Returning to Initial State (deleting this screenshot will undo the action and return to initial state)

### v2

@@ 5.2. Method
To input the UI elements from the UI action traces into the evaluation prompts, we detect UI elements from each screen (Zhang et al., 2021) and format them as an HTML string (Wang et al., 2022). For MLLMs (multimodal LLMs), we use the raw screenshot data. We used 5 different LLMs and MLLMs for the evaluation: (1) GPT4, (Text Only), (2) GPT4 (MLLM), (3) Gemini 1.5 Flash (Text Only), (4) MM1.5 (MLLM) (Zhang et al., 2024a), and (5) Ferret-UI (MLLM) (You et al., 2023), an MLLM trained on UI understanding tasks.
@@ 6. Discussion
In this section, we reflect on the research process and results of two studies. We discuss the limitations including: (1) the granularity of impact, (2) brainstorming vs. realistically performing actions, and promising future works, including different potential usages of the taxonomy.
@@ 6.2. Future Work
Individuals have different perceived feelings over “impacts.” Our taxonomy can act as a reasoning tool to guide AI agents in determining the level of impact and, ultimately, shaping how they respond to those impacts. By using this taxonomy, future AI agents could make their decision-making process more transparent and understandable, showing how they were calibrated to assess the effects of UI actions. Moreover, users could personalize their own policies based on these taxonomy categories. For instance, if an AI agent determines that reversing a UI action requires “multiple steps in a timely manner”, users could customize the outcome. They might set the decision to “minimum impact” if the action does not affect other users, or to “significant impact” if it changes how others perceive them. This potential of the taxonomy brings new possibilities to make AI agents’ decisions more accessible, customizable, and reliable by understanding different aspects of “impacts.”
Base for future UI agent research.
Our taxonomy lays the groundwork for understanding post-action impacts on UIs. As noted in the limitations, further research is needed to refine the granularity of impact assessments. Currently, our data annotations classify impacts into three categories: minimal impact (i.e., AI can proceed safely), moderate impact (i.e., AI should seek user confirmation), and significant impact (i.e., AI should halt and defer to human judgment). However, real-world scenarios are often more complex, with no one-size-fits-all policy for AI behavior. These policies are highly dependent on the specific domain and application. We encourage future researchers to build on our work by exploring the relationship between UI action impacts and the policies that AI agents should follow when navigating UIs. Another promising direction for future research is examining how UI design might adapt to better accommodate AI agents. For instance, can future UI designs minimize the need for user confirmations while maintaining effective collaboration with AI agents? Additionally, what new safety challenges might arise from such design changes?
@@ 7. Conclusion
This work aimed at understanding and evaluating the consequences of UI actions by introducing a comprehensive taxonomy of UI action impacts. Through collaborative workshops with domain experts and an extensive data synthesis study, we have captured a diverse range of UI actions that go beyond harmless browsing, focusing instead on actions with potential interactions with other parties that might need intervention. Our findings demonstrate that current datasets do not adequately represent the complexity of UI action interactions, particularly those with significant consequences. We show a significant gap in state-of-the-art LLMs’ ability to reliably understand the complexity of what happens as impact after UI actions.
@@ A.4. Adding Chain-of-Thought (CoT)
Roll Back Effects: Returning to Initial State (deleting this screenshot will und)